### v1

# Think Together and Work Better: Combining Humans’ and LLMs’ Think-Aloud Outcomes for Effective Text Evaluation

Abstract.
This study introduces InteractEval, a framework that integrates human expertise and Large Language Models (LLMs) using the Think-Aloud (TA) method to generate attributes for checklist-based text evaluation. By combining human flexibility and reasoning with LLM consistency, InteractEval outperforms traditional non-LLM-based and LLM-based baselines across four distinct dimensions, consisting of Coherence, Fluency, Consistency, and Relevance. The experiment also investigates the effectiveness of the TA method, showing that it promotes divergent thinking in both humans and LLMs, leading to the generation of a wider range of relevant attributes and enhance text evaluation performance. Comparative analysis reveals that humans excel at identifying attributes related to internal quality (Coherence and Fluency), but LLMs perform better at those attributes related to external alignment (Consistency and Relevance). Consequently, leveraging both humans and LLMs together produces the best evaluation outcomes. In other words, this study emphasizes the necessity of effectively combining humans and LLMs in an automated checklist-based text evaluation framework. The code is available at https://github.com/BBeeChu/InteractEval.git.
Large Language Model, Think-Aloud, Human-LLM Combination, Text Evaluation, Checklist
Figure 1. The overall process of prompting-based, checklist-based text evaluation models, and InteractEval (Ours)
1. Introduction
In recent years, the rapid advancement of large language models (LLMs) has led to a surge in research focused on text generation (Lee et al., 2022; Mysore et al., 2023; Schick et al., 2022; Yang et al., 2022; Lee et al., 2024a), not to mention the industries: Notion AI  [^1] https://www.notion.so/product, Jasper  [^2] https://www.jasper.ai/, Writesonic  [^3] https://writesonic.com/, sudoWrite  [^4] https://www.sudowrite.com/, and Cohesive  [^5] https://cohesive.so/, to name a few textual content generation solutions. Accordingly, the need for accurate mechanisms to evaluate the quality of automatically generated text has grown increasingly critical (Fu et al., 2023; Liu et al., 2023; Lee et al., 2024b; Wang et al., 2023), and led to various LLM-based evaluation approaches. LLM-based evaluation methods have introduced significant improvements by enabling reference-free evaluations using the chain-of-thought (CoT) prompting strategy (Liu et al., 2023) and adjusting scores based on output token probabilities (Fu et al., 2023). However, as shown in Figure 1 (A), these methods depend heavily on task-specific prompts and tend to only evaluate the general aspects of the summary, leading to reduced reliability. To address these challenges, checklist-based approaches (Lee et al., 2024b, a) introduce the use of checklists to evaluate text by breaking down rating dimensions into several key components. These methods attempt to enhance the accuracy of the evaluation by having the LLM focus on specific components when scoring and analyzing the summary, thereby assessing the summary based on detailed elements of the dimension.
Nevertheless, concerns persist regarding the manually created key components in these checklists. As shown in Figure 1 (B), the checklists are generated by LLMs through sub-questions derived from key components manually defined by the authors based on a rubric. This process raises two main concerns. First, the reliance on human-defined key components introduces a strong dependency on human judgment. According to previous studies, humans and models each possess distinct strengths (Pantcheva, 2023; Blattberg and Hoch, 1990; Wang et al., 2024a; Das et al., 2023). Thus, depending exclusively on human-defined key components could introduce bias into the evaluation process. Secondly, it raises the question of whether these components fully capture how the aspects intended to be evaluated in the rubrics’ rating dimensions manifest in the actual text. This approach follows a theory-based method (Plakans, 2013; Rodríguez, 2008), where a rating scale is designed based on established theories or pre-defined guidelines, such as rubrics. This method potentially overlooks the detailed features represented in example texts, focusing only on the rubric, and bring about a coarse-grained checklists (Brindley, 1998; Turner and Upshur, 2002).
This study introduces InteractEval, addressing the limitations of relying solely on either humans or machines by combining their ideas to mitigate biases that arise from over-reliance on one source. We merge the ideas of both LLMs and human experts derived from the think-aloud (TA) process, known for fostering divergent thinking (Ericsson and Simon, 1980; Ericsson, 6; Nielsen, 1994; Lewis, 1982), with rubrics and sample texts. These combined TA results are used to create unbiased and fine-grained outcomes for text evaluation (see Figure 1). We expect this synergy to complement humans and LLMs as they hold different strengths, and reduce bias and improve checklist quality (Liu et al., 2023; Wang et al., 2024a). Human experts excel in flexible, high-level reasoning (Wu et al., 2023; Blattberg and Hoch, 1990; Pantcheva, 2023), while models offer consistent output, even when processing a large amounts of information (Wang et al., 2024a; Das et al., 2023; Blattberg and Hoch, 1990). During the TA process, we merge performance-driven approaches, which review individual samples, with the theory-based methods. This makes humans and LLMs specifically consider how the aspects intended to be rated in a dimension are manifested in sample texts (Fulcher et al., 2011; Knoch, 2009; Jeffrey, 2015).
Our framework contains three main steps to combine the outcomes of TA process and create checklists based on them. First, after the TA by humans and LLMs, we collect thoughts from both humans and LLMs, referred to as text attributes, which are needed for rating specific dimensions of texts. Next, we prompt an LLM to extract key components from the merged ideas, to cluster the merged thoughts based on these components, and to generate questions related to each component. Lastly, we ask the LLM to create a checklist for rating a specific dimension of the given texts.
The research questions in this paper are:
  1. RQ 1.
    Does our text evaluation framework, which uses checklists designed with integrated attributes from human experts and LLMs through the TA method, outperform baseline text evaluation methods in terms of correlation with human judgments?
  2. RQ 2.
    Do the attributes generated through the TA method improve the evaluation performance of generated texts, based on the checklists constructed from those attributes?
  3. RQ 3.
    How effective is the combination of attributes from humans-TA and LLMs-TA for text evaluation using the checklists constructed from the TA-generated attributes?
2. Related Work
2.1. Text Evaluation with Large Language Models
LLM-based text evaluation has emerged to address the shortcomings of traditional text evaluation methods, such as BLUE (Papineni et al., 2002) or ROUGE-L (Lin, 2004), which highly depend on reference text (Fabbri et al., 2021; Gu et al., 2021; Guan et al., 2021). Research on text evaluation using LLMs (Liu et al., 2023; Fu et al., 2023; Kocmi and Federmann, 2023; Chan et al., 2023; Kim et al., 2024) has been actively progressing, starting with GPTScore (Fu et al., 2023) which score texts based on its token generation probability and G-Eval (Liu et al., 2023) which use the Chain-of-Thought (CoT) (Wei et al., 2022) technique to evaluate texts. Studies utilizing multi-agent collaboration (Chan et al., 2023; Kim et al., 2024) are particularly active, aiming to overcome the single-thread limitations (Wang et al., 2024b) of single-LLM approaches, which hinder scalability and speed when dealing with complex tasks like text evaluation, which requires taking multiple aspects of text quality into consideration. Despite these advantages, those methods are resource-intensive, incurring high costs in time and computation. They experience increased time to first token (TTFT) (Wang et al., 2024b), and still fall short in their robustness on text evaluation. CheckEval (Lee et al., 2024b) introduced a binary evaluation method using checklists to assess texts, based on the observation that LLMs excel at fine-grained tasks (Liu et al., 2022; Min et al., 2023) and are particularly effective for yes/no questions. However, the heavy reliance on humans (specifically the authors) in designing key components for checklist creation raises concerns about the reliability of the evaluation criteria (Lee et al., 2024b). Moreover, it still suffers from being coarse-grained by designing the key components solely based on the broad rubric definition of each dimension.
2.2. Combination of Human Intelligence and Artificial Intelligence
Numerous studies have demonstrated that AI decision-making, when combined with human interventions, outperforms decisions in isolation (Blattberg and Hoch, 1990; Wang et al., 2024a; Das et al., 2023; Revilla et al., 2023; Cohen et al., 2023). Historically, research has shown that human involvement improves the accuracy of statistical models in forecasting (Blattberg and Hoch, 1990; Clemen, 1989). Today, various studies continue to explore the effectiveness of human intervention in AI across different tasks. For example, in the retail industry, human involvement can significantly enhance an AI model’s predictions, especially when dealing with long time horizons and low uncertainty. Additionally, collaboration between humans and large language models (LLMs) improves the accuracy of data annotation (Wang et al., 2024a). LLM performance, notably, is highly dependent on specific tasks, datasets, and labels (Zhu et al., 2023; Ziems et al., 2024), and they often retain biases and errors in their outputs (Liu et al., 2023; Wang et al., 2024a), making human oversight crucial. Moreover, human-centered AI technologies remain essential for tasks like fact-checking (Das et al., 2023), highlighting the growing importance of human-AI interaction, particularly in high-level tasks like text evaluation within LLMs (Lee et al., 2024b).
Human experts and models each have unique strengths and weaknesses, making them complementary when used together. Humans are more flexible but can become fatigued or bored with large tasks, while models are more consistent and rigid (Blattberg and Hoch, 1990). Given the extensive prior knowledge (e.g. universal grammar (Montague et al., 1970; Chomsky et al., 2023)) humans possess, many studies in the human-in-the-loop (HITL) field aim to enhance the consistent performance of neural models by incorporating human expertise into data processing, model training, inference, and system development (Wu et al., 2022; Xin et al., 2018). Human elements, such as empirical knowledge and practical skills, help models learn and perform specific tasks, especially when dealing with sparse data (Wu et al., 2022; Kumar et al., 2019). Considering the biases in LLM performance (Liu et al., 2023; Wang et al., 2024a), human intervention can help mitigate these issues. Since most LLMs are trained with human feedback (Achiam et al., 2023; Touvron et al., 2023; Anthropic, 2024), and tasks like text generation and summarization can be improved with HITL (Ziegler et al., 2019; Stiennon et al., 2020), involving humans in checklist construction for text evaluation could enhance the quality of LLMs’ text ratings.
2.3. Think Aloud
The think aloud (TA) method, where participants verbalize their thoughts while performing experimental tasks, is a well-established technique in cognitive psychology and usability studies. Introduced by Ericsson and Simon (Ericsson and Simon, 1980), this method has been pivotal in understanding cognitive processes and identifying usability issues by capturing real-time user feedback (Nielsen, 1994; Lewis, 1982; Fan et al., 2019). There are two main types in the TA method: concurrent TA (CTA) and retrospective TA (RTA) (Ericsson, 6). In CTA, users verbalize their thoughts while performing a task, whereas in RTA, they recall and describe their thoughts and feelings after completing the task. Although both methods have their strengths, they also have limitations. To address them, a hybrid (HB) method has been developed that combines elements of both approaches (Alhadreti and Mayhew, 2018).
The TA method has been widely applied beyond usability studies, serving as a tool across various fields to develop new methodologies by capturing diverse ideas (Ericsson and Simon, 1980; Nielsen, 1994). For example, it was used to construct a rating scale for English writing assessment by having in-service teachers and experts think aloud their thoughts on relevant text attributes while reviewing sample essays and scoring dimensions (Wu, 2022). These insights were then consolidated into a detailed rubric, providing a structured approach to rubric creation. Building on (Wu, 2022), this study seeks to create a fine-grained rubric-based checklist for evaluating generated-texts. Using the TA method, we involve both humans and LLMs to identify specific text attributes for the rating dimensions-Coherence, Fluency, Consistency, and Relevance-ultimately refining and improving LLM-based evaluation.
3. Method
Figure 2. InteractEval Framework for Text Evaluation: (A) Think Aloud: Human experts and LLMs verbalize their thoughts to generate text attribute insights using sample texts and evaluation rubrics. (B) Checklist Construction: Insights are combined and categorized into key components, leading to the generation and validation of checklist questions. (C) Checklist-based Evaluation: The checklists are answered by an evaluator LLM to evaluate the summaries, with results aggregated into a final score.
3.1. Overview
InteractEval combines the TA outcomes from humans and LLMs to build checklists for reliable text evaluation. Inspired by the competitive and robust performance of CheckEval (Lee et al., 2024b) in text evaluation, this study adopts a checklist-based evaluation approach. InteractEval consists of three distinct stages, as illustrated in Figure 2. Initially, humans and LLMs participate in the TA process by independently brainstorming the text attributes to be considered when rating texts. These attributes are then merged and used to create checklists consisting of multiple boolean-style questions, requiring binary responses (Yes/No). The evaluator LLM reviews the text summaries based on the checklist and provides a response to each question. Finally, the ’Yes’ responses are tallied to generate a final score, which is then compared with the human-assigned ground truth scores. A key distinction from CheckEval is that InteractEval generates the components using the attributes generated by both human and LLM, reducing bias and including more fine-grained criteria (Pantcheva, 2023; Blattberg and Hoch, 1990; Brindley, 1998; Turner and Upshur, 2002).
3.2. Dimensions for Evaluation
The study uses SummEval (Fabbri et al., 2021) dataset, which includes a collection of news article summaries, for a case study. The dataset comprises source news articles on various topics, reference summaries, machine-generated summaries from different models, and human-evaluated quality scores (from 1.0 to 5.0) across four dimensions: Coherence, Fluency, Consistency, and Relevance. According to previous studies (Zhu and Bhat, 2020; Wang et al., 2024c), generated texts possess two key features: internal quality and external alignment. Internal quality refers to grammatical accuracy, readability, and structure, focusing on the intrinsic characteristics of the text. In contrast, external alignment evaluates adherence to prompts and factual accuracy, assessing how well the summary reflects and aligns with the source text. Based on the rubric from the original paper (Fabbri et al., 2021), we categorize Coherence and Fluency under internal quality, and Consistency and Relevance under external alignment. The detailed definition of each dimension is as follows:
  • Internal Quality
    • Coherence: The collective quality of all sentences. We align this dimension with the DUC quality question of structure and coherence whereby ”the summary should be well-structured and well-organized. The summary should not just be a heap of related information, but should build from sentence to a coherent body of information about a topic.
    • Fluency: The quality of the summary in terms of grammar, spelling, punctuation, word choice, and sentence structure.
  • External Alignment
    • Consistency: The factual alignment between the summary and the summarized source. A factually consistent summary contains only statements that are entailed by the source document. Annotators were also asked to penalize summaries that contained hallucinated facts.
    • Relevance: Selection of important content from the source. The summary should include only important information from the source document. Annotators were instructed to penalize summaries which contained redundancies and excess information.
3.3. Attributes Collection through Think Aloud
Figure 3. Human experts and LLMs are provided with dimension rubrics, sample texts for each dimension, and their corresponding scores. They then review the materials simultaneously, offering their thoughts regarding text attributes that should be considered when rating the dimension based on their knowledge.
In this stage, we collected the text attributes for constructing checklists of each dimension. For this, we utilized the TA method, which enhance divergent thinking (Ericsson, 6; Lewis, 1982). During the TA, participants verbalized their thoughts while conducting experimental tasks (Ericsson and Simon, 1980). Likewise, four human participants and four different LLM participants articulated the text attributes needed for rating the text quality of each dimension in this study. To minimize biases arising from individual members of each participant group, whether humans or LLMs, we included four members in each group. They did this while reading the definitions of each dimension along with four sample summaries per dimension as illustrated in Figure 3. As a result, the TA participants generated fine-grained attributes by considering the detailed features of each dimension as manifested in the example summaries. Between the CTA and the RTA (Ericsson, 6), which are two main methods of the TA, the current framework used the CTA, as it is straightforward to implement (Fan et al., 2019), avoids biases from post-task retrospection (Van den Haak et al., 2004; McDonald et al., 2012), and has been shown to have minimal impact on participants’ behavior (Hertzum et al., 2009).
3.3.1. TA by human experts
The human experts include two journalism and two linguistics specialists, chosen based on the nature of the news article summary dataset. We assumed that journalism experts would focus on aspects related to news content, while linguistics experts would concentrate on the language features of the texts. Their details are provided in Table 1. The experts conducted the process in real-time via online video calls, addressing each dimension—Coherence, Fluency, Consistency, and Relevance—separately. Specifically, we collected participants’ thoughts on each dimension through the following process. First, participants were provided with four news articles and summaries, actual scores corresponding to different score ranges, and the definitions of each dimension. Then, participants reviewed the news summaries corresponding to each score range and verbalized their reasoning for why they believe the news articles received the given scores. Finally, the participants’ verbalizations were recorded, transcribed, and the identified text attributes were compiled.
Expert Occupation Related Experience Language Use (English)
1 Ph.D Student Researching on digital and AI journalism Fluent (non-native)
2 Journalist News article generation with generative models Fluent (non-native)
3 University Lecturer Teaching English writing in a university Fluent (native)
4 University Lecturer Teaching English writing in a university Fluent (native)
Table 1. Information of human experts who participated in the TA process.
3.3.2. TA by LLMs
The LLMs include four different types of representative agent: GPT-4 (Achiam et al., 2023), Llama-3.1-8B-Instruct  [^6] https://llama.meta.com/, Claude-3.5-Sonnet  [^7] https://www.anthropic.com/news/claude-3-5-sonnet, and Gemini-1.5-Pro (Reid et al., 2024). The LLMs simulate a human-like TA process using the CTA method, processing text attributes needed for scoring each dimension in real-time. For this, we construct a prompt template with the identical materials given to humans during the TA: four news articles and summaries, corresponding scores, and dimension definitions. During the TA, we assign a persona (assessor) to the LLMs and make them read the materials and verbalize their thoughts on what to keep in mind when evaluating a targeting dimension. In addition, they are required not to be too much oriented to specific samples (See Appendix A for the prompt templates used in the LLMs’ TA).
3.4. Checklist Construction
In this stage, we merge the attributes collected from the previous TA process (Section 3.3) of both humans and LLMs. Then, we build the checklists using the combined text attributes. The checklist construction process includes four steps: 1) Component Extraction, 2) Attributes Clustering, 3) Question Generation, and 4) Question Validation. Each step can utilize an LLM, such as GPT series, and in our study, we use either GPT-3.5-turbo or GPT-4. The integration of attributes from both sources (humans and LLMs) is expected to create a synergistic effect, where the strengths of human expertise and LLMs complement each other, leading to a more refined and effective checklist (see Appendix B to F for the prompt templates used in each step). A key difference between InteractEval and CheckEval (Lee et al., 2024b) is how components are created. In CheckEval, humans (authors) manually derive components from a rubric. In contrast, InteractEval generates components from a combination of text attributes based on the TA outcomes of both humans (experts) and LLMs, using rubrics and sample texts. This approach reduces bias from relying on a single source and allows for more detailed checklists base on the fine-grained attributes (Pantcheva, 2023; Blattberg and Hoch, 1990; Brindley, 1998; Turner and Upshur, 2002).
3.4.1. Component extraction
The component extraction process identifies key themes related to a rating dimension from the text attributes generated in the prior TA process. For instance, a component related to coherence may involve attributes such as ”Structure”, ”Logical Flow” and ”Tone style”. This step narrows down the focus areas, laying the groundwork for generating key questions tied to these components. Using an LLM as an extractor, principle components are extracted by prioritizing the most pertinent attributes from both human experts and the LLMs. To avoid noise and redundancy, we prompt the extractor to eliminate ambiguous or repetitive components from the combined attributes. Also, we let the agent combine related attributes into cohesive entities. The number of components is limited to five or fewer to prevent excessive subdivision, which can lead to overlapping meanings and create noise. These components then guide the subsequent steps in checklist development.
3.4.2. Attributes clustering
After extracting the components from the combined group of attributes, an LLM clusters each attribute under the appropriate component to identify specific associated attributes. This clustering divides the dimension into segmented sections, leveraging the ideas of both humans and LLMs. Each section focuses on a specific aspect of the dimension, enabling more targeted evaluation. For example, under the ’Logical Flow’ component, attributes such as ”Evaluate whether the connections between different situations are well represented,” ”Check if the descriptions of the same subject are well-linked,” and ”Check if the sentences are related to each other” can be clustered.
3.4.3. Question generation
In this stage, the clustered attributes from the prior step are transformed into actionable ”Yes-No” questions aimed at probing the identified attributes. First, the LLM generates a key question for each component based on the corresponding attributes. For example, if there are five components, the LLM generates five key questions grounded on each component. Subsequently, the agent breaks down the key questions into detailed sub-questions to more precisely identify the relevant features within a summary. These questions form the core of the checklist, guiding the LLM evaluator in assessing the summaries.
3.4.4. Question validation
In the final stage, generated questions undergo rigorous review by an LLM for clarity, relevance, and effectiveness. The LLM is prompted to combine any redundant questions and drop irrelevant questions, ensuring the checklist is both reliable and practical for final use in rating texts. The questions should be answerable with ”Yes” or ”No” containing concepts of a rating dimension as well. After the validation process, the final questions compose the checklist which is used for text evaluation.
3.5. Checklist-based Evaluation
As mentioned before, we adopt checklist-based evaluating approach using the checklist created following aforementioned procedures. Therefore, an LLM evaluator is used to directly respond to each checklist question in a binary manner (Yes or No) during this stage. The number of positive responses (‘Yes’ answers) is then tallied. The proportion of positive responses to the total number of questions is scaled from 1.0 to 5.0 as the final score, as shown in Equation 1, to align with the score range of the ground truth scores in the dataset. This method assumes that the ratio of positive responses reflects overall text quality, with each question targeting a specific evaluation component. The scaled scores of each summary are compared with the ground truth scores assigned by humans.
(1) Final Score=Number of Positive ResponsesTotal Number of $\text{Final Score}=\frac{\text{Number of Positive Responses}}{\text{Total Number of Questions}}\times 5.0$
4. Experiment
The experiment consists of three phases aligned with the research questions. First, we compare our method, InteractEval, with baseline text evaluation methods. Next, we assess the effectiveness of the TA approach in creating checklists for text evaluation. Finally, we explore the impact of combining TA outcomes from humans and LLMs, comparing it to other TA conditions: Single-LLM-TA, Single-Human-TA, Multiple-LLMs-TA, and Multiple-Humans-TA. We employ both quantitative metrics, such as correlation with human evaluations, and qualitative analysis, focusing on the generated attributes.
4.1. Dataset
We evaluate our InteractEval framework using SummEval (Fabbri et al., 2021), a widely recognized benchmark in the text generation task particularly oriented to news article summarization. SummEval provides source texts, reference texts, and model-generated summaries, along with evaluation dimensions and human scores: Coherence, Fluency, Consistency, and Relevance. To assess our framework effectively and consider its generalizability, we sample 10% of the entire dataset, ensuring that the distribution of human annotations across each dimension and source text is uniformly represented in the sample. This is the way utilized in (Lee et al., 2024b) as well.
4.2. TA Condition Description
We examine the effectiveness of the TA process in generating attributes for the checklist used in text evaluation. We compare rating performances and qualitatively analyze the text attributes across five different TA conditions: Single-LLM-TA (a single LLM verbalizing its thought process, SL-TA), Single-Human-TA (involving one human expert, SH-TA), Multiple-LLMs-TA (with several LLMs, ML-TA), Multiple-Humans-TA (involving multiple human experts, MH-TA), and Combination-TA (combining human experts and LLMs, Comb-TA).
4.3. Quantitative Analysis
4.3.1. Measuring correlation
We measure the performance of InteractEval using sample-level correlation. This method computes correlation for each individual sample based on summaries from multiple source texts and then averages these correlations across all samples (Lee et al., 2024b). For each source text $x_{i}$, where $i\in\{1,2,\dotsc,I\}$, there are $M$ system-generated summaries $y_{i,m}$, with $m\in\{1,2,\dotsc,M\}$. Let $f_{\text{auto}}$ denotes the automatic evaluation metric (e.g., BLEU) and $f_{\text{human}}$ represents the human evaluation scores. The sample-level correlation $C^{\text{sample}}$ for each dimension is defined as:
(2) $C^{\text{sample}}_{(f_{\text{auto}},f_{\text{human}})}=\frac{1}{I}\sum_{i=1}^{I}g\Big{(}$ $[f_{\text{auto}}(y_{i,1}),\dotsc,f_{\text{auto}}(y_{i,M})],$
$[f_{\text{human}}(y_{i,1}),\dotsc,f_{\text{human}}(y_{i,M})]\Big{)},$
where $g$ is a correlation coefficient such as Spearman or Kendall’s tau.
4.3.2. Measuring difference
In this study, the Fisher’s transformation (Fisher, 1970) is employed as a measurement method to normalize and compare the correlation coefficients of the five separate TA conditions. Converting the correlation values into z-scores allows more accurate comparison and analysis of correlations, particularly in hypothesis testing and constructing confidence intervals.
4.3.3. Measuring similarity
We conduct text analysis to examine the similarity of generated attributes within the same dimension. This analysis evaluates whether each TA condition produces diverse attributes from both a lexical and semantic perspective. For lexical similarity, we use ROUGE (Lin, 2004) and Jaccard similarity (Jaccard, 1912; Niwattanakul et al., 2013), which focus on word overlap and sequences without considering meaning or context. For semantic similarity, we use cosine similarity, which measures semantic closeness rather than exact word matching.
4.3.4. Baselines
The following baselines, including widely used metrics (e.g. ROUGE-L (Lin, 2004), BLEU (Papineni et al., 2002), & METEOR (Banerjee and Lavie, 2005)), are compared with our InteractEval.
  • BERTScore (Zhang et al., 2019): measures text similarity using semantic embeddings generated by BERT (Devlin, 2018).
  • MoverScore (Zhao et al., 2019): improves BERTScore by integrating soft alignments and innovative aggregation methods for a more accurate similarity assessment.
  • BARTScore (Yuan et al., 2021): acts as a unified evaluator by utilizing the pretrained BART model (Lewis et al., 2019), which relies on the average likelihood of the model’s output.
  • UniEval (Zhong et al., 2022): functions as a unified multi-dimensional evaluator capable of evaluating different dimensions of text generation tasks.
  • G-Eval (Liu et al., 2023): employs LLMs for evaluating text quality, utilizing a chain-of-thoughts method and a form-filling framework to rate the quality of texts.
  • CheckEval (Lee et al., 2024b): uses checklists to assess text quality by breaking down evaluating dimensions and generating questions for each. Then, it gathers LLM responses to the checklist to calculate rating scores.
4.4. Qualitative analysis
We conducted a focused group interview (FGI) (Morgan, 1996; Dilshad and Latif, 2013) with four human experts (E1, E2, E3, and E4) to qualitatively analyze the attributes produced by humans (A) and LLMs (B). The experts compared both sets of attributes by responding to open-ended questions for each dimension (see Appendix G for the details of questions). These questions allowed us to gain deeper insights into the attributes and encouraged detailed expert feedback. Three of the experts had participated in the TA, while one had not. Initially, we planned for all four TA participants to conduct the qualitative analysis, assuming their familiarity would lead to more refined insights. However, one expert with linguistics expertise withdrew for a personal reason, so we invited another linguist, who had been also researching large language models, to participate. After collecting the experts’ responses through email, we identified keywords mentioned by the majority (at least three experts). We then filtered out answers containing the keywords. The results will be further discussed in section 5.3.2 and 5.3.3.
5. Results
5.1. Performance comparison with baselines (RQ1)
Internal Quality External Alignment Average
Models Coherence Fluency Consistency Relevance
$\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$
Non-LLM based
ROUGE-L 0.172 0.122 0.240 0.167 0.317 0.235 0.420 0.301 0.260 0.187
BLEU 0.028 0.019 -0.079 -0.057 0.036 0.030 0.423 0.308 0.023 0.017
METEOR 0.009 0.008 0.011 0.008 0.101 0.074 0.275 0.182 0.055 0.040
BERTScore -0.219 -0.151 -0.156 -0.113 -0.109 -0.076 0.462 0.476 -0.188 -0.134
MOVERScore 0.340 0.252 0.345 0.247 0.384 0.277 -0.434 -0.315 0.349 0.268
BARTScore 0.475 0.343 0.395 0.285 0.457 0.341 0.602 0.455 0.451 0.330
UniEval 0.567 0.411 0.605 0.447 0.613 0.476 0.640 0.476 0.555 0.412
LLM based (GPT-3.5-turbo)
G-Eval 0.412 0.393 0.354 0.333 0.338 0.322 0.361 0.321 0.366 0.342
CheckEval 0.440 0.347 0.393 0.327 0.600 0.515 0.305 0.252 0.436 0.360
InteractEval 0.583 0.536 0.630 0.604 0.734 0.702 0.614 0.576 0.640 0.605
LLM based (GPT-4)
G-Eval 0.619 0.470 0.629 0.482 0.664 0.517 0.617 0.464 0.632 0.483
CheckEval 0.573 0.428 0.632 0.493 0.706 0.611 0.570 0.438 0.620 0.493
InteractEval 0.650 0.603 0.798 0.768 0.783 0.760 0.626 0.587 0.668 0.640
Table 2. Sample-level correlations of different dimensions on SummEval benchmark dataset measured by Spearman ($\rho$) and Kendall tau ($\tau$). The best results are highlighted in bold, and the second-best results are underlined.
To evaluate the performance of InteractEval, we conduct a comparison against baseline evaluation methods. Building on previous studies (Fu et al., 2023; Fabbri et al., 2021; Lee et al., 2024b; Liu et al., 2023), we analyze the correlation between InteractEval and human annotation scores using Spearman and Kendall-tau coefficients to validate the reliability of our evaluations. For a fair comparison, we use the performance data of the non-LLM-based traditional method, G-Eval (GPT-4) and CheckEval as reported in (Lee et al., 2024b), since their experiments were conducted using the same experimental settings (sample-level correlation). We also reproduced the correlation results of G-Eval (GPT-3.5-Turbo) at the sample level using the source code provided by the original authors  [^8] https://github.com/nlpyang/geval.git, as none of previous work has included the results.
Table 2 presents the main results, classifying the four dimensions into two key features by which generated texts are typically assessed: internal quality and external alignment (Zhu and Bhat, 2020; Wang et al., 2024c). InteractEval with GPT-4 outperforms the traditional non-LLM-based metrics, such as ROUGE-L, BLEU, METEOR, BERTScore, MOVERScore, and BARTScore, in terms of correlation coefficients. Except for the Spearman correlation in the ”Relevance” dimension, InteractEval with GPT-4 generally outperforms UniEval. As noted in (Lee et al., 2024b), this suggests that using a checklist enables an LLM to evaluate generated text in a manner more aligned with human judgment than traditional automatic metrics.
When compared to G-Eval and CheckEval with the identical version of GPT, our model demonstrates superior performance. Moreover, even with GPT-3.5-Turbo, our framework outperforms other LLM-based baselines with GPT-4 or shows comparable performance. This indicates the cost-saving potential while ensuring reasonable performance of our proposed framework (Arefeen et al., 2024; Renze and Guven, 2024; Shekhar et al., 2024). These results highlight the broad capability of our proposed framework, showing that using TA and the checklist constructed through the combinatory TA of humans and LLMs lead to accurate and cost-effective evaluations across various text quality dimensions. Since InteractEval performs best when using GPT-4 as the evaluator, the remaining experiments were conducted with GPT-4.
5.2. Analysis of TA effectiveness (RQ2)
Internal Quality External Alignment Average
Models Coherence Fluency Consistency Relevance
$\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$
G-Eval 0.619 0.47 0.629 0.482 0.664 0.517 0.617 0.464 0.632 0.483
CheckEval 0.573 0.428 0.632 0.493 0.706 0.611 0.570 0.438 0.620 0.493
InteractEval (GPT-4) 0.620 0.598 0.633 0.600 0.738 0.724 0.550 0.530 0.635 0.613
InteractEval (Llama-3.1-8B-Instruct) 0.494 0.449 0.570 0.554 0.583 0.572 0.436 0.406 0.521 0.495
InteractEval (Gemini-1.5-Pro) 0.453 0.413 0.658 0.647 0.630 0.610 0.468 0.429 0.552 0.525
InteractEval (Claude-3.5-Sonnet) 0.632 0.581 0.467 0.427 0.522 0.462 0.523 0.470 0.536 0.485
Table 3. Sample-level correlations of different dimensions on the SummEval benchmark dataset were measured using Spearman ($\rho$) and Kendall-tau ($\tau$). Entire models utilized GPT-4 as the evaluator. The model names in parentheses in InteractEval refer to the models used for the TA process. The best results are highlighted in bold, and the second-best results are underlined, presented evaluator-wise.
The proposed framework initially collects various attributes from humans and LLMs that should be considered while designing checklists per dimension. Unlike previous methods, we apply TA to encourage LLMs to generate a wider range of text attributes, which form the basis for checklists across different evaluation dimensions. To examine how the TA process affects the evaluation performance of our framework, we compare its performance when using checklists constructed from a single LLM’s attributes under the TA process against other LLM-based baselines. The performance is shown in Table 3. As a result, the performance of InteractEval across four dimensions varies depending on the LLM used for the TA process. Still, our method with the single-LLM-TA outperforms other LLM-based baselines including the average, except in Spearman correlation for Relevance. Nevertheless, InteractEval using the outcomes of single-LLM TA with GPT-4 surpasses the baselines on average.
These results, shown in Table 3, suggest that even with just one LLM, the TA process leads to more refined checklists, improving evaluation outcomes. By encouraging divergent thinking (Ericsson and Simon, 1980; Ericsson, 6; Lewis, 1982), the TA process enables LLMs to capture more detailed aspects of the dimension rubrics based on sample texts. However, in the case of Relevance, our method did not outperform the baselines in terms of Spearman correlation, suggesting that the dimension may need further refinement in checklist design.
5.3. Comparative analysis of five TA conditions (RQ3)
As this study differentiates the TA conditions, we observe how performance varies based on differences in attributes generated through the think-aloud method under five conditions: SL-TA, SH-TA, ML-TA, MH-TA, and Comb-TA. We then delve deeper into the effectiveness of combining the outcomes of humans’ and LLMs’ TA for generating checklists. The SL-TA indicates the correlations averaged across four individual LLM conditions, and the SH-TA represents the performance averaged over four different expert conditions.
5.3.1. Performance comparison across different TA conditions.
Figure 4. Comparison of evaluation performance in Spearman correlations between single-TA and multiple-TA conditions across four dimensions.
Figure 5. Comparison of two correlations measures (Spearman & Kendall-tau) over four dimensions-Coherence, Fluency, Consistency, and Relevance. The bars represent the correlations of different TA conditions: SL-TA (Avg), SH-TA (Avg), ML-TA, MH-TA, and Comb-TA. The performances of SL-TA (Avg) and SH-TA (Avg) represent the average values across the four different checklists created by each LLM and human expert.
As shown in Figure 4, using the attributes generated by multiple humans or LLMs results in better evaluation outcomes, including improved average performance, compared to using the ones thought by a single human or LLM, though the difference is not statistically significant (Table 4). Moreover, Figure 5 demonstrates that combining the TA outcomes from humans and LLMs for checklist construction (Comb-TA) outperforms all single-TA (SL-TA, SH-TA) and multiple-TA (ML-TA, MH-TA) based checklists, with the exception of the Kendall-tau score for Coherence. However, the gap is negligible and does not significantly affect the overall average performance. The improvement is significant in Fluency and Consistency. This suggests that certain aspects of text quality benefit more from this combination than others.
(a) Comparison of evaluation performance. As Rouge-L (X-axis) decreases, Spearman correlation (Y-axis) increases. The red line shows the upward trend across all dimensions. The decreasing Rouge-L signifies an increasing diversity of attributes within each condition.
(b) Topic distribution and standard deviation analysis using LDA.
Figure 6. Analysis of attributes and evaluation performance.
Figure 6(a) shows the correlation between Spearman values and self-ROUGE-L scores for attributes within the same TA condition. This suggests that checklists based on diverse attributes (i.e., with low similarity) enhance evaluation performance, supporting the idea that the multiple-TA condition helps create more fine-grained checklists. Furthermore, combining human and LLM outcomes leads to even more fine-grained checklists, resulting in the highest performance and diversity.
To examine potential biases within the attributes of each TA condition, we conducted latent dirichlet allocation (LDA) (Blei et al., 2003), as shown in Figure 6(b). This method generated five topics based on the attributes from the Comb-TA condition. We then analyzed these topics to determine whether the attributes from other TA condition were concentrated in specific topics, providing insights into potential biases. The results show that attributes from single-TA conditions, whether human or LLM, tend to cluster into a few topics, with some topics receiving no attributes at all, indicating bias. In contrast, attributes from the multiple-TA and Comb-TA conditions are more evenly distributed across topics. Additionally, the Comb-TA condition has the lowest standard deviation in topic distribution, suggesting that combining the TA outcomes of multiple humans and LLMs results in a less biased checklist.
Internal Quality External Alignment Average
Methods Coherence Fluency Consistency Relevance
$\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$
L vs H SL-TA vs SH-TA 0.441 0.393 0.295 0.739 -0.914 -0.328 -0.324 0.937 -0.133 0.220
ML-TA vs MH-TA 0.565 0.866 0.563 0.277 -0.295 -0.409 -0.456 -0.455 0.057 0.228
S vs M SL-TA vs ML-TA 0.525 0.793 0.989 1.484 0.385 0.428 0.274 0.388 0.528 0.758
SH-TA vs MH-TA 0.648 1.266 1.257 1.022 1.005 1.166 0.142 -0.145 0.719 0.766
S vs C SL-TA vs Comb-TA 1.197 1.433 3.216** 3.54*** 2.967** 3.087** 1.318 1.602 1.241 1.659
SH-TA vs Comb-TA 0.755 1.040 2.921** 2.808** 3.881*** 3.415*** 1.643 1.523 1.374 1.439
M vs C ML-TA vs Comb-TA 0.672 0.640 2.226* 2.063* 2.581** 2.659** 1.044 1.213 0.713 0.901
MH-TA vs Comb-TA 0.106 -0.226 1.663 1.786 2.876** 2.249* 1.5 1.669 0.655 0.672
Table 4. Comparison of two correlations measures, Spearman ($\rho$) and Kendall-tau ($\tau$), over four dimensions-Coherence, Fluency, Consistency, and Relevance. The values are z-scores that indicate the statistical significance of the differences between the method pairs (L: LLM, H: Human, S: Single, M: Multiple, C: Combination), with asterisks (*) denoting the level of significance: * (p-vale <0.05), ** (p-value <0.01), and *** (p-value <0.001). The results are presented evaluator-wise.
Figure 7. Similarity of attributes in lexical and semantic perspectives. The bars represent each measurement of text similarity. The similarity of SL-TA (Avg) and SH-TA (Avg) represent the average values across the four different checklists created by each LLM and human expert. The lower the similarity, the more diverse the attributes are within a dimension.
5.3.2. Humans have strengths in internal quality
When comparing InteractEval’s performance using checklists based on LLM-only TA (SL-TA and ML-TA) with human-only TA (SH-TA and MH-TA) in both single and multiple settings according to Figure 5, human-only TA outperforms LLM-only TA in internal quality (Coherence and Fluency) of the generated texts. This suggests that humans excel in identifying attributes related to structure, semantic connections, and readability, which are critical to internal text quality. Their flexible thinking skills (Blattberg and Hoch, 1990) and innate understanding of universal grammar (Chomsky et al., 2023; Montague et al., 1970) allow them to generate higher-quality attributes that account for the hierarchical structure of texts (Pantcheva, 2023), a task that is more challenging for LLMs.
Qualitative analysis supports this conclusion. In the analysis, (A) refers to the attributes generated by humans and (B) refers to those generated by LLMs. For Coherence, humans created more quantitative and explicit attributes, making an evaluator easier to measure consistently. E3 commented, ”(A) offers simple, straightforward evaluations with easily quantifiable items, while (B) includes more qualitative, subjective items.” Additionally, humans focused more on logical order and organization. E1 noted, ”Attribute (A) emphasizes clear flow, meaningful connections, and sentence links for efficient information delivery, while (B) focuses more on comprehension but is less efficient.” For Fluency, humans provided attributes in more detail, while LLMs took a broader approach. LLM-generated attributes sometimes confused experts due to overlap with other dimensions. E1 remarked, ”Attribute (A) offers detailed and specific requirements, like including at least two interrogatives, while (B) provides more general guidelines.” E2 noted, ”Section (B) had many ambiguous items and concepts related to consistency.”
Additionally, the attribute similarities shown in Figure 7 support these findings. Human-generated attributes for internal quality are less lexically and semantically similar compared to LLM-generated ones, suggesting that humans’ TA produces more diverse attributes, allowing checklists to assess summaries across a wider range of features.
5.3.3. LLMs have strengths in external alignment.
In terms of external alignment (Consistency and Relevance), our framework using checklists based on LLM-only TA (SL-TA and ML-TA) shows superior performance-or minimal differences—compared to those based on human-only TA (SH-TA and MH-TA). This suggests that LLMs excel in identifying attributes related to external alignment, where both the source text and the summary must be considered simultaneously. The task requires processing large amounts of information, where LLMs’ vast context window and expansive memory excel (Pantcheva, 2023; Achiam et al., 2023; Wu et al., 2023). LLMs provide consistent attributes even when handling large data sets, adhering to their pre-trained policies (Blattberg and Hoch, 1990; Pantcheva, 2023).
Qualitative analysis supports these findings. For Consistency, LLMs explain the attributes in more detail and cover a broader scope than humans. E1 noted, ”Attribute (B) describes each point with more detail,” and E4 added, ”Attribute (B) fully explains more details compared to (A).” Additionally, LLMs provide systematic guidelines for detecting hallucinations, while human experts only provide attributes to check for false content. E2 commented, ”(B) gives detailed guidelines for identifying hallucinations, while (A) only shows how to detect them.” For Relevance, humans tend to create confusing attributes that could apply to multiple dimensions. One expert mentioned being confused, thinking that certain attributes were Coherence: ”In (A), some sentences could be interpreted as a coherence issue.” E4 also noted, ”Sometimes (A) felt like it belonged in another category.” LLMs, on the other hand, clearly reflect the rubric’s definition of Relevance and generate attributes covering a broader scope. Three out of four experts agreed on this. E3 said, ”(B) covers more areas and provides more accurate and detailed descriptions for each item according to the rubric.”
As a result, LLM-generated attributes lead to better checklist construction and improved evaluation performance in the dimensions related to external alignment, as shown in Figure 5. Moreover, Figure 7 shows that LLM-generated attributes for external alignment are less lexically and semantically similar compared to those created by humans, indicating that LLM-TA produces more diverse attributes, enabling checklists to assess summaries across a wider range of features.
6. Discussion
Figure 8. Comparison of evaluation performance averaged across four dimensions using two correlation metrics (Spearman and Kendall-tau). Single refers to the average performance of InteractEval with checklists based on SL-TA and SH-TA, Multiple refers to the average performance with checklists based on ML-TA and MH-TA, and Combination represents the performance using checklists based on Comb-TA.
6.1. Major Findings
The goal of this study is to develop a text evaluation framework, InteractEval, using checklists built from attributes generated through the combined TA processes of humans and LLMs. Our key findings are as follows. First, InteractEval outperforms both traditional non-LLM and LLM-based baselines in terms of correlation with human-assigned scores. Analysis of attribute topics and similarities shows that this improvement stems from reducing biases and enhancing fine-grainedness by incorporating both human and LLM perspectives, as well as rubrics and sample texts, supporting RQ 1.
Secondly, the TA process encourages divergent thinking in LLMs, generating a wider range of text attributes, which form the basis of the checklists. This is confirmed by the strong performance of checklists based on the attributes from single-LLM TA, supporting RQ 2.
Finally, in response to RQ 3, combining the TA outcomes of humans and LLMs mitigates biased attributes and increases their fine-grainedness, leading to more refined checklists and improved evaluation performance. Additionally, humans and LLMs exhibit distinct strengths in text evaluation. Humans excel in assessing internal text quality, where flexible thinking and an innate understanding of grammar are crucial. In contrast, LLMs are more effective at detecting external alignment, thanks to their ability to process large amounts of information consistently.
Therefore, combining the complementary outcomes of both humans’ and LLMs’ TA in text evaluation enhances overall performance, as shown in Figure 6(b), compared to using attributes from either humans or LLMs alone.
6.2. Implications
This study offers practical and research-related implications. As previously noted, text generation is rapidly expanding alongside advancements in LLMs. At the same time, the need for accurate methods to evaluate the quality of generated texts is gaining significant interest. InteractEval provides a cost-effective metric that shows reliable correlations with human assessments. Similarly, in research areas where reliable metrics for evaluating LLMs’ text generation abilities are scarce, our framework can serve as a useful testing method. Furthermore, in the context of rapidly increasing machine-centered automation, including LLMs, this study underscores the importance of human high-level thinking abilities and suggests a direction for effective human-machine interaction.
7. Limitations
In this study, we identified three main limitations. First, we tested our framework using the SummEval dataset (Fabbri et al., 2021) because it includes the two key features we focus on: internal quality and external alignment. As a result, our findings may not generalize to other datasets or text generation tasks, requiring further research. Secondly, during the experiment, we observed differences in the capabilities of humans and LLMs when constructing attributes for checklists. However, because comparing these capabilities was not the main focus of this study, we did not explore them in depth. Lastly, to ensure a fair comparison with baselines that primarily use GPT models, we did not use smaller LLMs, such as Llama-3.1-8B-Instruct, as evaluators. Future studies could include smaller LLMs to assess the proposed methods’ capabilities.
8. Conclusion and Future Work
Our work distinguishes itself from other LLM-based text evaluation studies. We introduce InteractEval, a framework that integrates human expertise and LLMs using the think-aloud (TA) method to generate attributes for checklist-based text evaluation. Our findings demonstrate that combining the strengths of humans and LLMs improves evaluation performance across the four dimensions, Coherence, Fluency, Consistency, and Relevance, highlighting the value of their collaboration (RQ1).
The study also emphasizes the effectiveness of TA in promoting divergent thinking in both humans and LLMs, helping to identify a broader range of attributes for each evaluation dimension (RQ2). This approach results in more comprehensive and refined checklists, ultimately enhancing the quality of text evaluation.
Additionally, we show that leveraging both human experts and LLMs together produces the best outcomes, reinforcing the importance of their combination (RQ3). Humans excel in identifying attributes related to internal quality (Coherence and Fluency), while LLMs outperform in external alignment (Consistency and Relevance).
Future research should assess the robustness and generalizability of this checklist-based approach, which is built on attributes generated by humans and LLMs through the TA process, using a variety of real-world text generation datasets. Additionally, exploring the strengths and weaknesses of humans and LLMs across different tasks could provide further insights. Ultimately, as the title suggests, this study demonstrates that when LLMs and humans think together, they can work better in text evaluation.
Appendix A Example of System Messages and Predefined Template for LLMs TA
A.1. System Message
You are an assessor who rate {Dimension} of the summary of a news article. You must read the summary samples and rubric for rating {Dimension}. While reading, write down your thoughts on what to keep in mind when evaluating the {Dimension} of the text, referring to the rubric. Avoid mentioning content specific to the samples. Instead, focus on general considerations for evaluating the {Dimension} of any summary based on the provided rubric.
A.2. Predefined Template
Rubric of {Dimension}:
{Rubric}
¡Sample 1¿
News Article: {Source Text}
Summary: {Summary}
Score: {Score}
¡End of Sample 1¿
¡Sample 4¿
News Article: {Source Text}
Summary: {Summary}
Score: {Score}
¡End of Sample 4¿
Please provide each consideration in a JSON format with the key of consideration index and the value of the consideration. Do not name each consideration. For example, {1: ”consideration 1”, 2: ”consideration 2”, 3: ”consideration 3”, 4: ”consideration 4”, 5: ”consideration 5”, 6: ”consideration 6”,…}
Appendix B Example of System Messages and Predefined Template for Component Extraction
B.1. System Message
Given a set of attributes, extract the components continually mentioned in the attributes based on the following conditions.
1. The components should be related to {Dimension} defined below.
2. Remove any components that are ambiguous or redundant.
3. The number of the components should not be more than 5.
Definition of {Dimension}: {Definition}
B.2. Predefined Template
Sentences: {Attributes}
Please, provide your answer in a List format with the list of components.
For example, [”component 1”, ”component 2”, ”component 3”, …]
Appendix C Example of System Messages and Predefined Template for Attributes Clustering
C.1. System Message
Given a set of attributes and components, group the attributes based on the components.
You can remove attributes that are redundant or confusing.
C.2. Predefined Template
Components: {Components}
Attributes: {Attributes}
Please, provide your answer in a JSON format with the keys of the components and the values of the list of attributes.
For example, {”component 1”: [”attribute 1”, ”attribute 2”], ”component 2”: [”attribute 3”, ”attribute 4”, ”attribute 5”], ”component 3”: [”attribute 6”], …}
Appendix D Example of System Messages and Predefined Template for Question Generation
The prompt template is inspired by the prompts introduced in (Lee et al., 2024b).
D.1. System Message
You are a key question generator for evaluating {Dimension}.
You have been provided with components along with their associated attributes. Your task is to:
1. Identify and extract common features that appear across the attributes of each component.
2. Based on these common features, construct a set of Yes-No questions related to each key component. These questions should be designed to effectively evaluate the accuracy and relevance of a summary of a news article.
Ensure that your response is well-organized, clear, and meets the highest standards of accuracy and reliability.
D.2. Predefined Template
The definition of {Dimension} and the set of components and their attributes corresponding to {Dimension} are provided below.
Use this information to generate Yes-No questions of the key components.
Each component and question must satisfy the following conditions:
1. Each component should have one corresponding question.
2. Each question must be answerable with ‘Yes’ or ‘No’.
3. Each question must incorporate concepts from the key component.
4. Each question should minimize the subjectivity of the rater’s judgment.
5. Formulate questions so that a ‘Yes’ answer is a positive response.
# Definition
{Rubric of a dimension}
# Components and attributes
{Components and attributes}
Please provide your answer in a JSON format with the keys of components and the values of questions.
For example, {”component 1”: ”question”, ”component 2”: ”question”, ”component 3”: ”question”, …}
The word ”component” should not appear in the keys of the answer format.
Appendix E Example of System Message and Predefined Template for Sub-Question Generation
The prompt template is inspired by the prompts introduced in (Lee et al., 2024b).
E.1. System Message
You are a sub-question generator. You must construct Yes-No sub-questions for each component.
E.2. Predefined Template
In this task, you need to create a question to evaluate the {Dimension} of the summary of the original document. The definition of {Dimension} and the questions corresponding to the key component of {Dimension} are provided below. Use them to generate sub-questions for each key question.
Each sub-question must satisfy the following conditions:
1. Each question must be answerable with ‘Yes’ or ‘No’.
2. Each sub-question must incorporate concepts from the key component.
3. Each question should minimize the subjectivity of the rater’s judgment.
4. The semantic redundancy between sub-questions should be minimized.
5. Formulate sub-questions so that a ‘Yes’ answer is a positive response.
6. Each sub-question must focus on asking about the presence of positive aspects rather than the absence of negative ones. For example, ask whether a summary avoids negative elements rather than contains them.
# Definition
{Rubric of a dimension}
# Components and corresponding questions
{Components and questions}
Please provide your answer in a JSON format with the key of components and the value of the list of sub-questions.
For example, {”component 1”: [”sub question 1”, ”sub question 2”,…] , ”component 2”: [”sub question 1”, ”sub question 2”, ”sub question 3”, …], ”component 3”: [”sub question 1”, ”sub question 2”, …], ”component 4”: [”sub question 1”, ”sub question 2”, ”sub question 3”, ”sub question 4”, …],…}
Appendix F Example of System Message and Predefined Template for Question Validation
F.1. System Message
You are a sub-question evaluator.
F.2. Predefined Template
In this task, you need to examine each sub-question for evaluating the {Dimension} of the summary of the original document. The definition of {Dimension} is provided below.
# Definition
{Rubric of a dimension}
Each sub-question must satisfy the following conditions:
1. Each question must be answerable with ‘Yes’ or ‘No’.
2. Each question must contain concepts of the {Dimension}.
3. Each question should minimize the subjectivity of the rater’s judgment.
4. The semantic redundancy between sub-questions should be minimized.
5. Formulate questions so that a ‘Yes’ answer is a positive answer.
6. Please combine any questions that ask about similar characteristics into a single question.
7. Each sub-question must focus on asking about the presence of positive aspects rather than the absence of negative ones. For example, ask whether a summary avoids negative elements rather than contains them.
# Sub-questions {Sub-questions}
Edit or exclude the problematic questions if it is necessary. Please provide your ultimate answer in a List format.
For example, [”sub question 1”, ”sub question 2”, ”sub question 3”, …]
Appendix G Open-ended questions for qualitative analysis
Please answer the following questions and analyze the differences between Attributes (A) and Attribute (B).
  • What are the key points measured in “Dimension” according to the given definition?
  • Please compare and analyze (A) and (B) with respect to the evaluation of the “Dimension.”
  • Please compare and analyze (A) and (B) with respect to fine-grainedness.
  • Please compare and analyze (A) and (B) regarding the diversity of their contents.
  • Please compare and analyze (A) and (B), focusing on aspects not covered in the questions above.
  • Please list the three most important differences between (A) and (B) overall. (You may repeat what has been already discussed.)